OpenStudy (johan14th):

(4x+f(x)) / (3−g(x)) can i get help to find the derivative using the quotient rule, thanks.

OpenStudy (johan14th):

@Loser66

OpenStudy (johnweldon1993):

We have to use the quotient rule? Alright well we know it can be written as \(\Large \frac{b(x)a'(x) - a(x)b'(x)}{b(x)^2}\) Knowing the top is a(x) and the bottom is b(x) \[\large a(x) = 4x + f(x)\] \[\large a'(x) = 4 + f'(x)\] \[\large b(x) = 3 - g(x)\] \[\large b'(x) = -g'(x)\] so \[\large \frac{(3 - g(x))(4 + f'(x)) - (4x + f(x))(-g'(x))}{(3 - g(x))(3 - g(x))}\] There is't really TOO much to simplify from there...but if you need to you can

OpenStudy (johan14th):

thank you! i have been stuck on this problem for a while now.